OpenStudy (princesssleelee):

After 3 months of living at his new house, Tony had 4 birds. Four months later, Tony still had 4 birds. Which of the graphs below models Tony's bird population? graph of line going through 4, 0 and 4, 8 graph of line going through 0, 4 and 7, 4 graph of line going through 1, 5 and 4, 8 graph of line going through 4, 0 and 7, 3

OpenStudy (princesssleelee):

@igigighjkl

OpenStudy (igigighjkl):

I'm here

OpenStudy (princesssleelee):

im thinking its the first answer

OpenStudy (igigighjkl):

Me too. :)

OpenStudy (princesssleelee):

I'll let you know if its correct.

OpenStudy (igigighjkl):

I'm thinking

OpenStudy (igigighjkl):

I think it might be the first one.

OpenStudy (igigighjkl):

(\/) (o.O) ^ ^ I made a bunny XD

OpenStudy (princesssleelee):

lmfaooooo!

OpenStudy (princesssleelee):

ok heres another! Given the function f(x) = The quantity of 4x minus 2, divided by 3, which of the below expressions is correct? f−1(x) = The quantity of 2 minus 4x, divided by 3. f−1(x) = The quantity of 3x minus 2, divided by 4. f−1(x) = The quantity of negative 4x minus 2, divided by 3. f−1(x) = The quantity of 3x plus 2, divided by 4.

OpenStudy (princesssleelee):

f^-1 (its an exponent) and the commas represent a fraction line.

OpenStudy (princesssleelee):

maybe the 4th once?

OpenStudy (igigighjkl):

What grade are you in?! o.O

OpenStudy (princesssleelee):

12th.

OpenStudy (igigighjkl):

I'M I SIXTH!!! XD

OpenStudy (princesssleelee):

wtfffff lmfao

OpenStudy (princesssleelee):

do you understand this question?

OpenStudy (igigighjkl):

not really, I'll call one of my fans.

OpenStudy (igigighjkl):

@pooja195

OpenStudy (princesssleelee):

lol alrighty

OpenStudy (igigighjkl):

@pooja195

OpenStudy (igigighjkl):

Sorry she couldn't help nor I.

OpenStudy (princesssleelee):

its alright im going to post another